OpenStudy (anonymous):

What is the equation of the line that passes through ( 1,5) and (4,11)? a. y = 6 b. y = x + 2 c. y = 2x + 3 d. y = 3x + 4 i wanna say the answer is B i still need help tho

OpenStudy (anonymous):

eq of line through two points is \[y-y1=\frac{ y2-y1 }{ x2-x1 }\left( x-x1 \right)\]

OpenStudy (anonymous):

so its not b . i really dont know how to solve this

OpenStudy (anonymous):

@surjithayer the equation is confusing

OpenStudy (anonymous):

\[y-1=\frac{ 11-5 }{4-1 }\left( x-1 \right)\] solve it.

OpenStudy (vjordillo):

C

OpenStudy (vjordillo):

@surjithayer, it should be (y-5)

OpenStudy (anonymous):

yes you are correct ,i typed wrong.thank you.